### v1

# Space-based mm/mg-scale Laser Interferometer for Quantum Gravity

Abstract
The experimental verification of the quantum nature of gravity represents a milestone in quantum gravity research. Recently, interest has grown for testing it via gravitationally induced entanglement (GIE). Here, we propose a space-based interferometer inspired by the LISA Pathfinder (LPF). Unlike the LPF, our design employs two smaller gold–platinum test masses, each weighing the milligram scale, surrounded by a shield below 2 K, and positioned side by side with a millimeter scale separation. This configuration enables the detection of GIE through simultaneous measurements of differential and common-mode motions. We simulate quantum measurements of these modes taking into account gas damping, black-body radiation, and cosmic-ray collisions to estimate the integration time for GIE detection. Our results show that GIE can be demonstrated with a few modifications to the LPF setup.
Introduction— Although extensive research on quantum gravity has been conducted, limited evidence has been obtained from astronomical observations [1, 2, 3, 4, 5, 6, 7]. This is primarily because the Planck scale, where quantum-gravity effects become significant, is far beyond the scales achievable in current experiments. Recently, novel methods have been proposed to bypass this difficulty and test whether Newtonian gravity exhibits quantum properties in the non-relativistic regime [8, 9]. These methods are based on a fundamental theorem in quantum information theory, which states that quantum entanglement cannot be generated through the LOCC (Local Operations and Classical Communication) [10]. Demonstrating that entanglement can be generated via gravity would provide direct evidence of its quantum nature.
Toward this end, several experiments have been proposed [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]; however, due to the weak nature of gravity, achieving this remains a challenge. There are three possible approaches to overcome this challenge: (1) signal amplification [19, 16, 18], (2) direct noise reduction [11, 12, 13, 14, 20, 17, 23], and (3) development of noise-tolerant measurement methods [15, 21, 22, 12]. The experimental setups can be categorized into two types: one targets massive objects, where gravitational interactions can dominate more easily, particularly in optomechanical systems [19, 21, 14, 17, 23], while the other focuses on microscopic systems [11, 12, 13, 20, 12], where quantum control is more easily achievable. A system combining these two components is also proposed [16, 15].
Here, we propose a cavity-free laser interferometer to optimize the second approach in optomechanical systems, in contrast to previous cavity-based studies [19, 21, 14, 17, 23]. Our proposed experiment closely resembles the space-based interferometer LISA Pathfinder (LPF) [24], which marked the beginning of a new era in interferometric experiments; however, it differs in the following four key aspects: (1) The test masses are reduced to the milligram scale to suppress decoherence caused by blackbody radiation; (2) The test masses are placed close to enable the generation of gravitationally induced entanglement (GIE); (3) Both differential and common-mode motions are measured to enable the detection of GIE via Kalman filtering; and (4) The masses are surrounded by a shield below 2 K, and both modes of motion are further cooled via feedback using a high-pass filter. Using the space environment, the resonance frequency of the test mass can be reduced below 1 mHz free from ground vibrations, making it possible to observe the GIE within the total integration time of about 200 days.
Figure 1: Experimental setup. The differential and common-mode of test mass 1 (TM1) and test mass 2 (TM2) are measured by laser interferometry with a wavelength of $1064\ \mathrm{nm}$.
Theory— We consider a system where two mirrors (TM1 and TM2) of equal mass $m$ are coupled via gravity. Two laser beams are directed at the test masses as shown in Fig. 1. The results of the homodyne measurement are
$Y_{\pm}=\bm{C}_{\pm}\bm{x}_{\pm}+y_{\pm}^{\rm in},$ (1)
where $Y_{\pm}$ represent the phase quadratures, $\bm{C}_{\pm}=(-\alpha_{\pm},0,0)$ denote the optomechanical coupling constants, and $\bm{x}_{\pm}=(q_{\pm},p_{\pm},r_{\pm})^{T}$ denote the position, momentum, and auxiliary variables used to describe feedback cooling (for details see Appendix A). The terms $y^{\rm in}_{\pm}$ represent vacuum (white) noise, with variances given by $\langle(y_{\pm}^{\rm in})^{2}\rangle=1$. In addition, we introduce vacuum noise $x^{\rm in}_{\pm}$ for the amplitude quadratures. Throughout this work, the subscripts $+$ and -- refer to the common mode and the differential mode, respectively. The optomechanical coupling constants $\alpha_{\pm}$ are given by $\sqrt{16\omega_{c}P_{\rm in}/(m\Omega_{\pm}c^{2}})$, where $\omega_{c}$ is the laser frequency and $P_{\rm in}$ the incident laser power, $m$ the mass of each mirror, $\Omega_{+}=\Omega$ and $\Omega_{-}=\Omega\sqrt{1-\delta}$ the resonance frequency of the common and differential modes. Here, $\delta=4Gm/(\Omega^{2}L^{3})$ denotes the gravitational coupling between the two mirrors, the resonance frequency for each mirror is denoted as $\Omega$, and $L$ represents the average distance between the mirrors. In this paper, we assume $m/L^{3}=\rho\Lambda$, where $\rho$ is the mass density of the mirror and $\Lambda$ is the factor determined by the shape and configuration of the mirrors [21].
The motion of the mirrors is given by
$\dot{\bm{x}}_{\pm}=\bm{A}_{\pm}\bm{x}_{\pm}+\bm{w}_{\pm}.$ (2)
Here, $\bm{A}_{\pm}=\begin{bmatrix}0&\Omega_{\pm}&0\\ -\Omega_{\pm}-g_{0}^{\pm}\omega_{\rm fb}^{\pm}&-\Gamma&-g_{0}^{\pm}\omega_{\rm fb}^{\pm}\\ -\omega_{\rm fb}^{\pm}&0&-\omega_{\rm fb}^{\pm}\end{bmatrix}$ parameterize the state-space model, where $\Gamma$ is the mechanical dissipation rate, $\omega_{\rm fb}^{\pm}$ are the cutoff frequency of the highpass filter and $g_{0}^{\pm}$ are the feedback gain. The noise terms are given by $\bm{w}_{\pm}=\left(0,\sqrt{2\Gamma}p^{\rm in}_{\pm}-\alpha_{\pm}x^{\rm in}_{\pm}+\frac{g_{0}^{\pm}\omega_{\rm fb}^{\pm}}{\alpha_{\pm}}y^{\rm in}_{\pm},\frac{\omega_{\rm fb}^{\pm}}{\alpha_{\pm}}y^{\rm in}_{\pm}\right)^{T}$, where ${p}^{\rm in}_{\pm}$ represent thermal (white) noise with variances $\langle({p}^{\rm in}_{\pm})^{2}\rangle=2n_{\pm}+1$, and $n_{\pm}$ are the average phonon occupation number.
In a linear quantum measurement described by Eqs. (1) and (2), the signals $Y_{\pm}$ inherently include shot noise, denoted by ${y}_{\pm}^{\rm in}$, while the mirror positions $q_{\pm}$ are intrinsically perturbed by radiation pressure noise, represented by $\alpha_{\pm}{x}_{\pm}^{\rm in}$. Based on the results of the measurements, the Kalman filter produces the optimal estimate $\hat{\bm{x}}$, which can be used to analyze the GIE, as follows [25, 26]:
$\dot{\hat{\bm{x}}}_{\pm}$ $=\bm{A}_{\pm}\hat{\bm{x}}_{\pm}+\bm{K}_{\pm}(Y_{\pm}-\bm{C}_{\pm}\hat{\bm{x}}_{\pm})$ (3)
$\dot{\bm{V}}_{\pm}$ $=\bm{A}_{\pm}\bm{V}_{\pm}+\bm{V}_{\pm}\bm{A}_{\pm}^{T}+\bm{N}_{\pm}$
$-(\bm{V}_{\pm}\bm{C}_{\pm}^{T}+\bm{L}_{\pm})(\bm{V}_{\pm}\bm{C}_{\pm}^{T}+\bm{L}_{\pm})^{T},$ (4)
where $\bm{K}_{\pm}=\bm{V}_{\pm}\bm{C}_{\pm}^{T}+\bm{L}_{\pm}$ are the Kalman gain and $\bm{V}_{\pm}=\langle(\bm{x}_{\pm}-\bm{\hat{x}}_{\pm})(\bm{x}_{\pm}-\bm{\hat{x}}_{\pm})^{T}\rangle$ are the conditional covariance. In addition, $\bm{N}_{\pm}=\langle\bm{w}_{\pm}\bm{w}_{\pm}^{T}\rangle$ and $\bm{L}_{\pm}=\langle y^{\rm in}_{\pm}\bm{w}_{\pm}\rangle$. Eqs (3) and (4) compute state estimate $\hat{\bm{x}}_{\pm}$ using a recursive predict-correct cycle. First, the Riccati equation (4) determines the theoretical conditional covariance $\bm{V}_{\pm}$ and the optimal gain $\bm{K}_{\pm}$, a calculation carried out independently of the measurement data. This predetermined gain is then used to fuse measurements to correct the state prediction, yielding minimum-variance estimate.
The Gaussian state entanglement can be characterized by the entanglement negativity defined by
$E_{N}=-\frac{1}{2}\log_{2}\left[\frac{\Sigma-\sqrt{\Sigma^{2}-4\text{det}\bm{\mathcal{V}}}}{2}\right].$ (5)
Here, $\bm{\mathcal{V}}$ is the conditional covariance matrix of the individual test masses (see Appendix B) and $\Sigma=\det\bm{\mathcal{V}}_{1}+\det\bm{\mathcal{V}}_{2}-2\det\bm{\mathcal{V}}_{12}$, where $\bm{\mathcal{V}}_{1}$, $\bm{\mathcal{V}}_{2}$ are the covariance matrices of TM1 and TM2 normalized by the frequency $\Omega$, respectively. $\bm{\mathcal{V}}_{12}$ represents the gravity-induced correlation matrix between individual masses. According to the separability condition for two-mode Gaussian states, the systems are entangled if and only if $E_{N}>0$ [27]. To verify whether quantum entanglement is present from Eq. (5), it is sufficient to confirm $\Sigma-\text{det}\bm{\mathcal{V}}>1$ for $\Sigma>2$.
The colored region in Fig. 2 shows the region where GIE is generated. Its generation rate must be faster than the thermal decoherence rate, as shown below:
$2\Gamma(2n_{+}+1)<\delta\Omega.$ (6)
From the above inequality, the crucial condition for the GIE can be written as
${\left({\Gamma T/2\pi\over 10^{-18}{\ \rm Hz\cdot K}}\right)}{\left({\rho\over 20{\ \rm g/cm^{3}}}\right)}^{-1}\left({\Lambda\over 2}\right)^{-1}<2.4,$
where $T$ represents the temperature of the test masses. The required $\Gamma T/(2\pi)$ is about $10^{-18}\,\mathrm{Hz\cdot K}$.
Figure 2: Contour plot of $E_{N}$ in the $\alpha_{+}$-$\Gamma T/(2\pi)$ plane. The green dashed line indicates the threshold for generating the GIE in Eq. (6), the red curve shows the boundary where the relaxation time due to the Kalman filter is equal to half the thermal decoherence time (see Appendix A). The star symbol corresponds to the parameters used in the simulation.
Sources of decoherence— As inherently unavoidable sources of decoherence, we consider gas damping, black-body radiation, and cosmic-ray collisions. First, the dissipation rate caused by gas, involving particles with mass $m_{\rm atom}$, pressure $P$, and temperature $T_{\rm env}$ is given by [28]
${\Gamma_{\rm gas}}$ $=$ ${PR^{2}\sqrt{8\pi m_{\rm atom}}\over m\sqrt{k_{B}T_{\rm env}}}{}\left(1+{h\over 2R}+{\pi\over 4}\right),$ (8)
where $k_{B}$ is the Boltzmann constant, and $h$ and $R$ denote the height and radius of the mirror, respectively.
Second, following [29], the total decay rate from thermal blackbody photons is estimated as $\Gamma_{\rm scat}+\Gamma_{\rm abs}+\Gamma_{\rm em}\,$where we define
$\Gamma_{\rm scat}$ $={8!{8}\hbar R^{6}\over 9\pi m}\left({k_{B}T_{\rm env}\over\hbar c}\right)^{8}\zeta(9){\rm Re}\left[{\varepsilon-1\over\varepsilon+2}\right]^{2}$ (9)
for the scattering by thermal photons, and
$\Gamma_{\rm abs}$ $={16\pi^{5}\hbar R^{3}\over 189m}\,\left({k_{B}T_{\rm env}\over\hbar c}\right)^{5}{\rm Im}\!\Biggl{[}\frac{\varepsilon-1}{\varepsilon+2}\Biggr{]}$ (10)
for absorption. Here, $\hbar$ is the reduced Planck constant, $\varepsilon$ is the dielectric constant, and $\zeta(z)$ is the Zeta function. For emission, $\Gamma_{\rm em}$ is obtained by replacing $T_{\rm env}$ in Eq. (10) with the temperature of the test masses.
Finally, the acceleration noise due to cosmic-ray collisions is calculated using Geant4 [30] and a cosmic ray flux model [31]. By removing the effects of large events caused by cosmic-ray collisions through data processing and control, the required shielding thickness to sufficiently reduce the noise can be limited to about 1 m for 18 mg test masses. The duty cycle — defined as the fraction of time during which the data is usable for analysis — is about 56% (for details see Appendix C).
Simulation— The parameter $\Sigma-\det(\bm{\mathcal{V}})$ depends on the fourth power of the variance components and therefore does not follow a Gaussian distribution. Consequently, instead of using error propagation, we estimate the integration time required to demonstrate GIE through simulations. First, the sample paths of the differential and common mode motions (hereafter referred to as the ’true values’), subject to thermal noise, radiation pressure noise, and feedback noise, are computed using the Euler–Maruyama method [32] with a sampling rate of 10 Hz. Second, the measurement data $Y_{\pm}$ are generated by multiplying these paths by the optomechanical coupling constants $\alpha_{\pm}$ and adding shot noise. Finally, the optimal state estimates $\hat{\bm{x}}$ are obtained by applying the Kalman filter, defined in Eqs. (3) and (4), to the simulated measurement data $Y_{\pm}$.
In the simulation, we set $\Omega_{+}/(2\pi)=7.7\times 10^{-4}\ \mathrm{Hz}$, $\alpha_{+}=0.081\ \mathrm{Hz^{1/2}}$, $\Gamma T/(2\pi)=10^{-18}\ \mathrm{Hz\cdot K}$, $\rho=21\ \mathrm{g/cm^{3}}$, $g_{0}^{\pm}=0.2$, and $\omega_{\rm fb}^{\pm}=5\Omega_{\pm}$, corresponding to the star symbol in Fig. 2. From these parameters, we obtain $\delta=0.48$ and $\Sigma-\det(\bm{\mathcal{V}})=1.02$. Since the parameters are configured to keep the feedback noise sufficiently low, we can extract the shot noise level, the combined thermal and radiation pressure noise level, and the effective susceptibility by fitting the measurement data to the following expression:
$S_{\rm fit}^{\pm}(\omega)=S^{\pm}_{\rm tot}|\chi^{\pm}_{\rm eff}(\omega)|^{2}+S^{\pm}_{\rm bg},$ (11)
where $S^{\pm}_{\rm bg}$ are the spectral background floor, $S^{\pm}_{\rm tot}$ the sum of force noise, $\chi^{\pm}_{\rm eff}=\Omega_{\pm}/(\omega^{2}-\Omega_{\pm}^{2}+i\omega\Gamma+g_{0}^{\pm}\omega_{\rm fb}^{\pm}\Omega_{\pm}i\omega/(i\omega+\omega_{\rm fb}^{\pm}))$ the effective mechanical susceptibility modified by the feedback. Using these fitted parameters, we numerically solve the Lyapunov equation - given by the first line of Eq. (4) - to compute the unconditional variances, $\bm{V}_{\pm}^{\rm un}=\langle\bm{x}_{\pm}\bm{x}_{\pm}^{T}\rangle$. The difference between the unconditional variances and the estimated variances $\hat{\bm{V}}_{\pm}=\langle\hat{\bm{x}}_{\pm}\hat{\bm{x}}_{\pm}^{T}\rangle$ yields the conditional variances $\bm{V}_{\pm}$, according to Eve’s law [33].
Figure 3: An example of the differential-mode ASD, $S_{q_{-}}^{1/2}$, measured over $10^{7}$ seconds. Blue shows the true values, light blue the estimate, black the measured data. Orange is shot noise, green is radiation pressure noise, magenta is thermal noise, yellow is feedback noise, and red circles show fit results. The inset shows the time-series data $q_{-}$ from 0 to 2000 seconds, with colors matching the main figure. Due to shot noise, the measured signal has an amplitude of about 200, causing the inset’s background to appear completely black.
An example of the differential mode amplitude spectral density (ASD) is shown in Fig. 3. The ASD is obtained using Welch’s method [34], with 50% overlap and a Hanning window. From about 100 simulation runs, we compute the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$, $\bm{V}_{+}$, and $\bm{V}_{-}$, with four integration times: $10^{5}$ s, $10^{6}$ s, $5\times 10^{6}$ s , and $10^{7}$ s (for details see Appendix D).
Results and Discussion— Let the theoretical value of $\Sigma-\det(\bm{\mathcal{V}})$ be denoted by $\beta$, and define $\Delta=\beta-1.$ Then the probability of observing GIE can be given by:
$P_{\mathrm{obs}}^{\mathrm{GIE}}=\int_{1}^{1+2\Delta}p\bigl{(}\ \Sigma-\det(\bm{\mathcal{V}})\bigr{)}\,\mathrm{d}\bigl{(}\Sigma-\det(\bm{\mathcal{V}})\bigr{)}$ (12)
where $p(\Sigma-\det(\bm{\mathcal{V}}))$ is the probability density function. For each measurement time, the cumulative probabilities $P_{\text{obs}}^{\text{GIE}}$ are 8.1%, 17.9%, 55.6%, and 72.7%, respectively. As the result demonstrates an improvement proportional to the square root of the measurement time, extrapolating it suggests that the measurement time required to achieve a probability of 99.7% is $1.8\times 10^{7}$ s. Although the measurement time becomes 372 days when the 56% duty cycle is taken into account for 18 mg test masses, it is the same as the one-year operational period of LPF.
For comparison, we calculate that the values of $P_{\text{obs}}^{\text{GIE}}$ are 0%, 3.3%, 15.9%, and 25.9%, respectively, for $\delta=0.28$ (which corresponds to $\Omega_{+}/(2\pi)=0.001$ Hz). In this case, the value of $\Sigma-\det(\bm{\mathcal{V}})$ is 1.007, and compared to the value of 1.02 at $\delta=0.48$, there is a three-fold difference in $\Delta$. Therefore, the integration time required to achieve 99.7% confidence is expected to differ by a factor of 10. The result supports this prediction.
Figure 4: Contour plot of $\log(\Gamma T/(2\pi))$ in the $\log m$-$T_{\rm env}$ plane at $\alpha_{+}=0.081\ \mathrm{Hz^{1/2}}$.
In terms of decoherence, the red region in Fig. 4 shows the region where the total dissipation given by Eqs. (8), (9), and (10) satisfies $\Gamma T/(2\pi)\leq 10^{-18}\ \mathrm{Hz\cdot K}$. Here, we consider a cylindrical mass with 3% absorption at a wavelength of 1064 nm [35, 36], having radius $R=L\sqrt{2/\pi}$ and height $h=L$, which yields $\Lambda=2$ for $m/L^{3}=\rho\Lambda$ [21]. These test masses are surrounded by a shield at temperature $T_{\rm env}$. We assume that the residual gas is an ideal gas consisting of hydrogen at a pressure of $5\times 10^{-16}\ \mathrm{Pa}$ at 4 K. The test masses, with an emissivity of 0.04 [37], are cooled by radiation.
By increasing the mass of the test masses, the gas damping can be mitigated. However, this comes at the cost of increased decoherence because of blackbody radiation emission. This trade-off is important in vacuum, where the levitated mass can only dissipate heat via radiation. As the mass increases, the incident laser power must be increased to maintain the value of $\alpha_{\pm}$, causing radiative cooling to become less effective. Consequently, the equilibrium temperature increases. For large masses, blackbody radiation becomes more influential, whereas for small masses, gas damping becomes more significant. As a result, the optimal mass scale for the GIE is on the order of milligrams. For example, when a 100 mg test mass is placed inside a shield cooled to 2 K and cooled by radiation, it attains an equilibrium temperature of 11 K under an absorbed laser heating power of 0.6 nW. This results in a value of $\Gamma T/(2\pi)=9.3\times 10^{-19}\ \mathrm{Hz\cdot K}$.
In principle, two key conditions must be validated on Earth before realizing the proposed space-based experiment. First, to observe the GIE, the gravitational interaction between the test masses must dominate over all other forces. Since gravitational coupling between milligram-scale objects has already been demonstrated [38], this condition can be satisfied. Second, achieving extremely high vacuum ($<10^{-15}$ Pa) and cryogenic temperatures ($\sim$2 K) in the space environment requires precise engineering design and quantitative validation. Although the required vacuum level is challenging, the ground-based experiment has already achieved $5\times 10^{-15}$ Pa at 4.2 K by cryopomping [39]. For the space mission, a 1 K–class Joule–Thomson cryocooler [40]—with a lifetime exceeding 3 years and a nominal cooling power of 10 mW at 1.7 K—is suitable. A cryocooler capable of cooling down to 0.05 K has also been developed [41]. The minimum value of $\Gamma T/(2\pi)$ is $2.6\times 10^{-19}\ \mathrm{Hz\cdot K}$, achieved at $T_{\mathrm{env}}=0.05\ \mathrm{K}$ and $m=18\ \mathrm{mg}$, as marked by the × in Fig. 4. Since the saturated vapor pressure decreases exponentially as the temperature drops [42, 43], it is preferable to lower the temperature as much as possible. Furthermore, key technical noise sources such as drag-free control, magnetic field effects, and charge-induced noise have already been characterized on LPF [24, 44, 45, 46]. This technological continuity indicates that the challenges faced by the proposed experiment can be reduced by applying systems already demonstrated in space.
Summary— Experimental investigation of the quantum nature of gravity is crucial to advancing modern physics, yet it remains a challenge. In this study, we demonstrate that such an investigation is marginally achievable with current technology. To this end, we simulated quantum measurements of two adjacent mg-scale test masses separated by a mm-scale distance, using a cavity-free interferometer. By improving the vacuum level less than $10^{-15}$ Pa, surrounding the masses with a shield below 2 K, and reducing the resonance frequency $\Omega/(2\pi)$ to $7.7\times 10^{-4}$ Hz, GIE can be demonstrated.
Acknowledgement— We thank Markus Arndt and Ichiro Arakawa for valuable discussions. We thank Koichi Ichimura and Tomoki Fujii for providing the Geant4 source codes. N.M. is supported by JST FORESTO Grant No. JPMJFR202X. K.Y. and A.M. are supported by JSPS KAKENHI Grant No. JP23H01175, and D.M. is supported by JSPS Overseas Research Fellowships.
Appendix A State space model
Here, we present the detailed derivation of Eqs. (1) and (2). These equations can also be applied to a dark-fringe Michelson interferometer with test masses in both arms.
First, let us denote the complex amplitude of the laser incident at $q_{1}=0$ by $\bar{a}+a^{\text{in}}$, where $\bar{a}$ is the mean value and $a^{\text{in}}$ is the vacuum fluctuation. We normalize it so that $|\bar{a}|^{2}$ equals the mean photon flux $P_{\text{in}}/(\hbar\omega_{c})$, and without loss of generality, we take $\bar{a}$ to be real: $\bar{a}=\sqrt{P_{\text{in}}/(\hbar\omega_{c})}$. The laser amplitude reflected from TM1 at a general position $q_{1}$ is, when referenced to $q_{1}=0$,
$a$ $=-\,\bar{a}\,e^{2ikq_{\text{zpf}}q_{1}}\;+\;a^{\text{in}},$
$\simeq-\,\bar{a}\;-\;2ikq_{\text{zpf}}\,\bar{a}\,q_{1}\;+\;a^{\text{in}},$ (A1)
where $k=\omega_{c}/c$ and $q_{\text{zpf}}=\sqrt{\hbar/(2m\Omega)}$. The phase quadrature of the light reflected from TM1, given by $Y_{1}=-\,i\,(a-a^{*})$, is
$Y_{1}$ = $=-\,4kq_{\text{zpf}}\,\bar{a}\,q_{1}\;+\;y^{\text{in}}$
$=-\,(\alpha/\sqrt{2}\,)\,q_{1}+y^{\text{in}},$ (A2)
where $\alpha=\sqrt{16\omega_{c}P_{\text{in}}/(m\Omega c^{2})}$. Then, the laser is reflected off TM2, either on the same side or on the opposite side, depending on whether the laser is used to detect the common mode or the differential mode, respectively. After reflection, the phase quadrature is
$Y_{\pm}=\mp\,(\alpha/\sqrt{2}\,)\,q_{2}+Y_{1},$ (A3)
where the upper and lower sign correspond to the common and differential modes, respectively. Thus, we recover the observation equation Eq. (1), with the following definition of the mechanical variables of the two modes:
$q_{\pm}=\sqrt{\frac{\Omega_{\pm}}{\Omega}}\frac{q_{1}\pm q_{2}}{\sqrt{2}},~{}~{}~{}~{}p_{\pm}=\sqrt{\frac{\Omega}{\Omega_{\pm}}}\frac{p_{1}\pm p_{2}}{\sqrt{2}}.$ (A4)
Second, to derive the equation of motion, Eq. (2), we first consider the case without the feedback cooling. In this case, the equation of motion for the test masses are given by
$\dot{q}_{i}=$ $\Omega\,p_{i}~{}~{}(i=1,2),$ (A5)
$\dot{p}_{1}=$ $-\Omega\,q_{1}-\Gamma\,p_{1}+\frac{\delta}{2}(q_{1}-q_{2})$
$+\sqrt{2\Gamma}\,p^{\text{in}}_{1}-\frac{\alpha}{\sqrt{2}}(x^{\text{in}}_{+}+x^{\text{in}}_{-}),$ (A6)
$\dot{p}_{2}=$ $-\Omega\,q_{2}-\Gamma\,p_{2}-\frac{\delta}{2}(q_{1}-q_{2})$
$+\sqrt{2\Gamma}\,p^{\text{in}}_{2}-\frac{\alpha}{\sqrt{2}}(x^{\text{in}}_{+}-x^{\text{in}}_{-}).$ (A7)
Here, $\alpha x^{\text{in}}_{\pm}$ represent the radiation pressure noise. Although they acquire time delay during the propagation between the test masses, it is far smaller than the time scale of oscillator and negligible. By combining the above equations along with Eq. (A4), we obtain
$\dot{q}_{\pm}=$ $\Omega_{\pm}\,p_{\pm},$ (A8)
$\dot{p}_{\pm}=$ $-\Omega_{\pm}\,q_{\pm}-\Gamma\,p_{\pm}+\sqrt{2\Gamma}\,p^{\text{in}}_{\pm}-\alpha_{\pm}\,x^{\text{in}}_{\pm}.$ (A9)
Next, we take feedback cooling into account, which motivates us to introduce the auxiliary variable $r(t)$ as follows. For notational simplicity, we omit the subscripts denoting the modes $\pm$ here.
The dynamics under feedback control is described by the above equations of motion with an additional force term $F_{\mathrm{fb}}$ on the right-hand side of Eq. (A9). For feedback cooling, this term is chosen as
$F_{\mathrm{fb}}=-\,g_{0}\,\Omega\,\hat{p},$ (A10)
where $g_{0}$ is the dimensionless feedback gain, so that cooling adds an effective damping $g_{0}\,\Omega$. We estimate $\hat{p}(t)$ from the phase-quadrature record $Y(t)$ via a first-order high-pass filter:
[…]
using $\dot{Y}/(-\alpha\Omega)\approx p$ and where $\omega_{\mathrm{fb}}$ sets the estimator’s bandwidth. Adding Eq. (A11) with Eq. (A10) to (A9) yields a non-Markov system with colored noise. To restore Markovianity, we define
[…]
whose evolution is
$\dot{r}$ $=-\,\omega_{\mathrm{fb}}\,r+\frac{\omega_{\mathrm{fb}}}{\alpha}\,Y$
$=-\,\omega_{\mathrm{fb}}\,r-\omega_{\mathrm{fb}}\,q+\frac{\omega_{\mathrm{fb}}}{\alpha}\,y^{\mathrm{in}}.$ (A13)
Together, Eqs. (A9), (A10), (A12), and this $\dot{r}$–equation form the Markovian system given in Eq. (2). In the $(q,p,r)$ formulation the feedback noise is treated exactly. If one attempted cooling by increasing $\Gamma$ in a model with $(q,p)$, the feedback noise would be omitted, which would violate the uncertainty relation for the conditional covariance. For a general linear system whose operators are governed by the equation of motion [Eq. (2)] and the observation equation [Eq. (1)], the expectation values and covariances of the conditional state are described by the quantum Kalman filter given by Eqs. (3) and (4).
Finally, to gain further insight, we present an analytical solution in an effective theory that neglects feedback noise. In the steady state, each element of the conditional covariance matrix is given by the following expressions [47]:
$V^{\pm}_{qq}=\displaystyle{\gamma_{\pm}-\Gamma\over\alpha_{\pm}^{2}}$ (A14)
$V^{\pm}_{qp}=\displaystyle{(\gamma_{\pm}-\Gamma)^{2}\over 2\alpha_{\pm}^{2}\Omega_{\pm}}$ (A15)
$V^{\pm}_{pp}=\displaystyle{(\gamma_{\pm}-\Gamma)(2\Omega_{\pm}^{2}+\gamma_{\pm}^{2}-\Gamma\gamma_{\pm})\over 2\alpha_{\pm}^{2}\Omega_{\pm}^{2}}.$ (A16)
Here, we introduce the characteristic frequency scales $\gamma_{\pm}$ of the Kalman filter, which are given by
$\gamma_{\pm}=\sqrt{\Gamma^{2}-2\Omega_{\pm}^{2}+2\Omega_{\pm}\sqrt{\Omega_{\pm}^{2}+(2\Gamma(2n_{\pm}+1)+\alpha_{\pm}^{2})\alpha_{\pm}^{2}}}.$
Considering the susceptibility of the position including the filter, it is expressed, just as in the original $\Gamma$, by
$\chi(\omega)$ $\propto$ $\frac{1}{\Omega_{\pm}^{2}-\omega^{2}-i\,\omega\,\gamma_{\pm}\ }\,.$ (A18)
Thus, $\gamma_{\pm}$ represent the full width at half-maximum (FWHM). To demonstrate the GIE, the condition $\gamma_{+}/2>n_{+}\Gamma$ represents an approximate threshold, as depicted by the red curve in Fig. 2 of the main text.
Appendix B Covariance matrix
The covariance matrix for individual test masses is extracted from the conditional covariances $V_{\pm}$ as follows. Among the components of $V_{\pm}$ representing correlations of $(q_{\pm},p_{\pm},r_{\pm})$, those involving $r_{\pm}$ are redundant, and the information of the two modes consists essentially of the following.
$\bm{\tilde{V}}_{\pm}=\left(\begin{array}[]{cc}V^{\pm}_{qq}&V^{\pm}_{qp}\\ V^{\pm}_{qp}&V^{\pm}_{pp}\end{array}\right).$ (B21)
Then, the covariance of the total system $\bm{\mathcal{V}}$, based on the two entangled test masses, is given by [17]
$\bm{\mathcal{V}}\equiv\Bigl{(}\begin{array}[]{cc}\bm{\mathcal{V}}_{1}&\bm{\mathcal{V}}_{12}\\ \bm{\mathcal{V}}_{12}&\bm{\mathcal{V}}_{2}\end{array}\Bigr{)}=\bm{S}\Bigl{(}\begin{array}[]{cc}\bm{\tilde{V}}_{+}&0\\ 0&\bm{\tilde{V}}_{-}\end{array}\Bigr{)}\bm{S}^{\text{T}},~{}~{}$ (B26)
$\bm{S}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cccc}1&0&1/(1-\delta)^{1/4}&0\\ 0&1&0&(1-\delta)^{1/4}\\ 1&0&-1/(1-\delta)^{1/4}&0\\ 0&1&0&-(1-\delta)^{1/4}\end{array}\right),$ (B31)
where $\bm{\mathcal{V}}_{1}$, $\bm{\mathcal{V}}_{2}$ are the covariance matrices of test masses 1 and 2 normalized by frequency $\Omega$, respectively. $\bm{\mathcal{V}}_{12}$ represents the gravity-induced correlation matrix between the test masses and $\bm{S}$ is the operation of the beam splitter.
Appendix C Cosmic-ray collisions
To evaluate shielding performance against cosmic-ray-induced test mass excitations, we performe Monte Carlo simulations using Geant4 [30]. The test mass is modeled as a cylindrical Pt-Au alloy (27% Pt and 73% Au by mass), located in the center of a spherical multilayer shield. We consider a mass value of 18 mg for the test mass. The test mass is modeled with radius $R=L\sqrt{2/\pi}$ and height $h=L$, which yields $\Lambda=2$ for $m/L^{3}=\rho\Lambda$.
[…]
We record the momentum kicks delivered to the test mass by cosmic-ray collisions. To realize GIE, the number of phonons generated by collisions must remain below at most one, throughout the GIE generation time $2\pi/(\delta\Omega)$. In practice, if a phonon excitation does occur, measurements are paused until the test mass has returned to its initial state. By increasing the feedback cooling gain $g_{0}$ to unity after each excitation, the effective quality factor $Q$ is reduced to 1. Consequently, the relaxation time scale to the initial state is approximately $2\pi/\Omega$, and the initialization time is weighted according to the noise magnitude. By performing the simulation described above, we find that the duty cycle, the fraction of time the measurement process is active within a given period, reaches 56% with 1.26 m of iron shielding. The average total number of large events during a measurement period of $2\pi/(\delta\Omega)$ is 0.895. The top panel of Fig. 5 shows the simulation result.
Figure 5: Simulation of test mass vibrations caused by cosmic-ray collisions, for a mass of 18 mg and $\delta=0.48$. Top panel: The time series data for the first $10^{5}$ seconds. The red segments (on a gray background) indicate regions excluded by data processing. The black dots indicate collision events with cosmic rays. Bottom panel: The averaged power spectral density obtained exclusively from the blue segments that were not excluded during data processing and whose durations exceed $2\times 2\pi/(\delta\Omega)$. It is comparable to the thermal noise level (dashed red).
In terms of the acceleration noise level due to cosmic-ray collisions, the required value for the amplitude spectral density (ASD) is calculated by the fluctuation-dissipation theorem [48]:
$\sqrt{S_{a}}$ $\simeq\frac{3.1\times 10^{-18}\ {\rm m}}{\rm{s}^{2}\sqrt{\text{Hz}}}\sqrt{\left(\frac{18~{}\text{mg}}{\rm{m}}\right)\left(\frac{\Gamma T/2\pi}{10^{-18}~{}\text{Hz}\cdot\text{K}}\right)}.$ (C32)
The requirement is met with 1.26 meters of iron shielding, as shown in the bottom panel of Fig. 5, where the acceleration noise is normalized by the zero-point amplitude. Although a total thickness of 1.5 m may not be impossible, it should be made thinner. It can be achieved through optimization of the shielding materials, implementation of active shielding with superconducting coils, and reduction of the test mass size.
Finally, in terms of electron charging, the rate due to secondary electrons generated by primary cosmic-ray protons is $7\times 10^{-5}\,\mathrm{e/s}$ (2200 electrons per year), which is approximately six orders of magnitude lower than the value measured by LPF [45]. The ratio of the gravitational potential energy $|\phi_{\rm Grav.}|=Gm^{2}/L$ and the Coulomb potential energy $|\phi_{\rm Coul.}|=N^{2}e^{2}/(4\pi\varepsilon_{0}L)$ is
$\biggl{|}{\phi_{\rm Grav.}\over\phi_{\rm Coul.}}\biggr{|}=3\times 10^{9}\left({m\over 100\ {\rm mg}}\right)^{2}{1\over N^{2}},$ (C33)
where $N$ is the number of electrons. This means that the effect of the Coulomb potential is negligible as long as $N<5\times 10^{4}$ for $m=100$ milligram objects.
Appendix D Probability density functions
Figure 6: The upper and middle panels display the statistical distributions of $\Sigma-\det(\bm{\mathcal{V}})$, $\bm{V}_{+}$, and $\bm{V}_{-}$. The lower shows $P_{\mathrm{obs}}^{\mathrm{GIE}}$ as a function of measurement time.
We compute the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$, $\bm{V}_{+}$, and $\bm{V}_{-}$, with four integration (measurement) times: $10^{5}$ s, $10^{6}$ s, $5\times 10^{6}$ s , and $10^{7}$ s. In the top and middle panels of Fig. 6, we show the probability density functions of $\Sigma-\det(\bm{\mathcal{V}})$ and of $\bm{V}_{+}$ and $\bm{V}_{-}$, respectively, for three integration times: $10^{6}$ s (blue), $5\times 10^{6}$ s (green), and $10^{7}$ s (red). The asterisks indicate the respective theoretical values. Compared to the differential mode, the conditional variance of the common mode exhibits a slight deviation from the theoretical value; this asymmetry arises because the laser power is equal in both modes, and the higher resonance frequency of the common mode makes it more susceptible to the influence of shot noise. In the lower panel of Fig. 6, circle and star symbols represent $P_{\mathrm{obs}}^{\mathrm{GIE}}$ for $\delta=0.28$ and $\delta=0.48$, respectively. The blue and red lines represent the cases where the cumulative probability increases in proportion to the square root of the integration time.